A rectangular room has walls 6 m and 10m long and 3m high. The sidewalls have absorptivity ?=0.05. The floor has a carpet with ?=0.45, and the roof is made of wood with ?=0.15. A machine is placed in the room and gives a background noise level of 65dB re 20
Answer #1

02_01_12.img_assist_custom-290x128.jpg

VOLUME=LBH

=6*10*3

=180m3

area=2(60+30+18)

=216m2

=0.161*180/216*0.65

=0.2064seconds

dB=10log(I/I0)

from this find I

=